Use the Taylor series to find the values of ln 0.4 accurate to 10^-3 . Use the integral remainder.


1
Expert's answer
2021-10-27T11:03:27-0400

Solution:

Using Taylor series expansion:

"\ln(1+x)=x-\dfrac{x^2}{2}+\dfrac{x^3}{3}-\dfrac{x^4}{4}+\dfrac{x^5}{5}-..."

Put "x=-0.6"

"\ln(1-0.6)=(-0.6)-\dfrac{(-0.6)^2}{2}+\dfrac{(-0.6)^3}{3}-\dfrac{(-0.6)^4}{4}+\dfrac{(-0.6)^5}{5}-...
\\ \Rightarrow \ln 0.4\approx-0.6-\dfrac{0.36}{2}-\dfrac{0.216}{3}-\dfrac{0.1296}{4}-\dfrac{0.07776}{5}
\\=-0.899952"

"\approx -0.900" [nearest to "10^{-3}" or 0.001]